# Unsupervised Multimodal Graph-based Model for Geo-social Analysis

### v1

@@ 1 Introduction
Social media platforms continuously generate vast amounts of data, including posts and comments, likes, shares, images, and videos. This content is rich, diverse, and real-time, offering a valuable resource for understanding how people interact, express opinions, form communities, and respond to events. The widespread adoption of smartphones with built-in location-tracking capabilities has added another dimension to this landscape: geographic context. By enabling users to share their locations, these devices have facilitated the integration of geospatial data into online content, enhancing the contextual depth of social media analysis. A growing body of research underscores the importance of geo-location data in social media research. Prior studies have demonstrated its utility across a wide range of applications, from disaster response [goodchild2010crowdsourcing] to public health monitoring [Signorini2011]. In particular, geo-referenced content from micro-blogging platforms has been leveraged to model and monitor phenomena such as earthquakes [Havas2021Portability], floods [Huang2018Flood], refugee movements [Havas2021Refugee], and disease outbreaks [Stolerman2023Disease].
Previous social media research has often been focusing on semantic topic modeling, a fundamental analytical algorithmic technique to uncover latent thematic structures within large volumes of user-generated text. A range of scholarly works have utilized topic modeling to reveal latent thematic structures within such texts [Resch2018] [Hannigan2019]. However, the resulting topic representations are rarely contextualized with additional data modalities. Early efforts to analyze geo-social textual media in a multimodal setting typically employed sequential workflows, wherein each modality was processed in isolation through distinct modules [yin2011geographical].
While these methods have demonstrated utility in various applications, their modular and linear design often hinders the ability to capture complex interdependencies between modalities. Furthermore, this sequential structure can introduce dependencies between processing stages, rendering the overall output sensitive to the order in which analyses are performed. More recent approaches introduced advancements toward integrating multiple modalities such as time, geographic space, semantic topics, and sentiments in the analysis of social media [gonzalez2024neural] [hanny2025]. However, they still depend on numerous isolated components and fragmented intermediate processing steps, falling short of delivering a truly end-to-end processing methodology. Nonetheless, inherent heterogeneity of input modalities in this case has already added to the complexity of the task. For example, text is symbolic and sequential, while location data is numeric and geospatial, often resulting in misaligned feature representations and inconsistencies during fusion. The lack of a unified machine learning model constrains cross-modal interactions and introduces fragile interdependencies, whereby variations in preprocessing or processing order can significantly impact both performance and interpretability.
This paper presents a new methodology for multimodal geospatial machine learning, that jointly embeds semantic and geospatial information in a unified graph-based space, using end-to-end learning with a composite loss to generate clusters that are both semantically coherent and spatially interpretable. The proposed methodology is application-agnostic and can, without loss of generality, be adapted to diverse forms of multimodal data integration, making it well-suited for a wide range of data fusion scenarios and downstream tasks across domains. To capture cross-modal relationships in a holistic and structured manner, we construct a unified multimodal embedding space based on a graph structure, where we can adapt diverse data types, including text, geolocation, and optionally emotions or temporal signals, into a shared representation space. In this formulation, nodes in the graph represent user-generated post embeddings, while edges encode relational signals such as semantic similarity, geographic proximity, or user relations. This enables the model to attend to multimodal interactions in a context-aware fashion.
At the core of our methodology, we train an artificial neural network in an unsupervised manner to aggregate information from a node’s neighbors, where a modified message passing mechanism allows scaling by the corresponding edge weights. A key innovation of our model is the use of a composite loss function inspired by reference-less clustering evaluation approaches. To enable robust and discriminative representation learning, the loss integrates contrastive, coherence, and alignment objectives. The contrastive component preserves local semantic structure by pulling similar nodes closer to each other in the embedding space while pushing dissimilar ones apart. The coherence objective encourages the formation of well-defined clusters by maximizing intra-cluster similarity and inter-cluster separation. At the same time, the alignment term stabilizes training by ensuring each node remains close to the centroid of its assigned cluster, thus enhancing representational consistency. By embedding these objectives directly into the learning process, the model is able to discover clusters that are both semantically meaningful and spatially coherent. Our approach is specifically tailored to support disaster management efforts. In emergencies triggered by natural hazards such as floods, geo-social media provides rich multimodal data from affected communities, enhancing situational awareness for first responders. Through a case study of the 2017 Hurricane Harvey in the USA, we demonstrate how geographically fine-grained multimodal clusters of social media posts closely align with the disaster’s observed impacts, illustrating the practical value of our method in real-world crisis scenarios. The framework itself is designed to be application-agnostic and, without loss of generality, can be adapted to various forms of multimodal data integration. It is well-suited for diverse data fusion scenarios and downstream tasks across different domains. Empirical evaluations demonstrate that our approach outperforms existing baselines in topic coherence, spatial compactness, and interpretability, providing a scalable and principled framework for multimodal social media analysis.
In summary, this work makes the following distinct contributions:
@@
  • Introducing a multimodal graph-based framework that seamlessly integrates multiple modalities (specifically semantic and geospatial data) into a unified representation space.
  • Proposing a domain-agnostic methodology for multimodal data analysis, underpinned by two unsupervised architectures, MonoGraph and MultiGraph, that adapt to diverse data characteristics and application needs, positioning the approach as a versatile framework for research and practice in fields such as disaster management, public health, and social media analytics.
@@
  • Developing a composite loss function that combines contrastive, coherence, and alignment objectives, ensuring the generation of semantically coherent and spatially interpretable clusters.
  • Demonstrating the effectiveness of our approach through a real-world disaster case study, where geographically fine-grained multimodal clusters of social media posts closely align with observed impacts.
@@ 2.1 Topic Modeling
Topic modeling is a widely used technique for discovering latent semantic topics within large document collections, serving as a variant of automated content analysis that assigns one or more “topic codes” to each document, thereby providing a structured thematic summary [hoyle2022content]. Among the foundational methods, Latent Dirichlet Allocation (LDA) [blei2003latent] introduced a fully generative probabilistic framework in which each document is modeled as a mixture over latent topics, and each topic is represented as a multinomial distribution over words. The generative idea of LDA has been extended to incorporate additional dimensions such as sentiment analysis [Lin2009Joint], temporal dynamics [Blei2006Dynamic], and social media user information [zhao2011comparing], enriching the interpretability and applicability of topic models across domains. However, LDA and its variants fundamentally rely on the bag-of-words assumption, which ignores word order and semantic relationships between words, limiting their effectiveness in handling short, noisy, or highly contextual texts such as tweets and social media posts [hong2010empirical]. To overcome these limitations, embedding-based topic modeling techniques have recently emerged. Dieng et al. [dieng2020topic] introduced the Embedded Topic Model (ETM), which integrates pre-trained word and sentence embeddings with LDA’s generative framework to better capture semantic relationships. Sia et al. [sia2020neural] further demonstrated that clustering high-dimensional semantic embedding vectors, derived from neural language models, can serve as a powerful alternative strategy, achieving performance comparable to traditional topic models. Building on these embedding-based approaches, Angelov [angelov2020top2vec] proposed Top2Vec, which leverages document embeddings and clustering algorithms to identify topics without requiring the number of topics to be pre-specified. This method utilizes techniques like Uniform Manifold Approximation and Projection (UMAP) and Hierarchical Density-Based Spatial Clustering of Applications with Noise (HDBSCAN) for dimensionality reduction and flexible clustering. Grootendorst [grootendorst2022bertopic] further advanced this line of work with BERTopic, which incorporates transformer-based contextual embeddings (e.g. from BERT) combined with refined clustering strategies, consistently outperforming traditional LDA and Top2Vec in both coherence and interpretability. Despite their progress, these models generally follow a sequential and fragmented framework and often struggle to balance semantic accuracy, interpretability, and scalability, particularly when applied to the diverse and noisy nature of social media data.
@@ 2.2 Spatio-semantic Topic Modeling
As location-aware social media platforms proliferate, spatially grounded topic modeling has gained increasing attention. This line of work focuses on uncovering geographically coherent semantic patterns by integrating geospatial signals into the semantic topic modeling process. Early efforts extended classical models like LDA with spatial priors. For example, Spatial LDA (sLDA) [Yuan2013Who] and Geo-aware Topic Models [Hong2012Discovering] include location-based information as latent variables or observed metadata, enabling the discovery of region-specific topics. Similar models have been applied to a variety of domains, such as understanding natural disasters [Resch2018], flood mapping [Huang2018Flood], and epidemic monitoring [Stolerman2023Disease]. However, most spatial topic models treat location and text independently or as loosely coupled variables, which often leads to weak cross-modal alignment. In addition, spatial topic models are frequently built upon modular architectures, making them difficult to scale or generalize. More recent multimodal approaches attempt to combine geolocation, text, and other modalities, often through hybrid processing pipelines [gonzalez2024neural] [hanny2025]. While these methods capture richer context, they extract each modality feature in isolation and then merge features post hoc, limiting their ability to jointly reason across modalities. Furthermore, lack of unified data space, and the representational gap between e.g. symbolic (textual) and continuous (geospatial) data introduces alignment challenges and interpretability issues in downstream clustering or topic interpretation.
@@ 2.3 Graph-based Topic Modeling
Several studies explored graph-based approaches to better capture semantic relationships and improve topic coherence. For example, Wang et al. [wang2015community] constructed word co-occurrence graphs from corpora and applied community detection algorithms, such as the Louvain method, to identify topical clusters, demonstrating superior performance on short text datasets. Similarly, Zhao et al. [zhao2018graph] proposed GSTM, a graph-based short text topic model that utilizes word co-occurrence graphs and community detection to improve topic coherence on sparse data. Beyond purely graph-based clustering, Graph Topic Models (GTM) have integrated Graph Neural Networks (GNNs) with topic modeling to jointly learn from document-word heterogeneous graphs, capturing both network structure and semantic information [zhou2020]. However, despite their ability to incorporate relational structure, GTMs still struggles to fully capture deeper or long-range semantic dependencies, limiting their effectiveness in modeling more abstract or nuanced topics. Earlier work by Newman and Girvan [newman2006finding] established a theoretical connection between community detection in bipartite document-word networks and topic model inference, highlighting the natural synergy between network science and probabilistic topic discovery. Extensions of LDA incorporating graph structures (such as Graph-LDA) have also been proposed to embed word similarity relations directly within the probabilistic framework, yielding more semantically coherent topics [doshi2015]. More recently, Graph2Topic [graph2topic2023] presented a topic modeling framework that leverags graph clustering methods on embedding-based sentence representations. This approach improves the quality and interpretability of topics without requiring a predefined number of topics, demonstrating the practical benefits of graph-embedding integration in topic discovery. Overall, graph-based topic modeling represents a promising direction that exploits the structural properties of the input data, allowing to integrate advances in network science and deep learning to overcome limitations of traditional models.
@@ 3 Methodology
Our approach leverages a set of geotagged social media posts, each represented by a semantic embedding $h_{i}$ (derived from the textual content) and a geographic coordinate $l_{i}$. To obtain rich semantic representations, we use Sentence-BERT (SBERT), which fine-tunes transformer-based language models in a Siamese architecture. This structure enables the model to directly optimize for sentence-level similarity, and has been shown in follow-up studies to effectively capture semantic proximity in short, informal texts such as tweets [Reimers2019]. To facilitate joint reasoning over these heterogeneous data types, we convert all modalities into a unified graph structure, serving as a common representation space. Graphs offer a flexible and expressive framework for capturing both local and global relationships within and across modalities. This design enables the neural model to propagate information through neighborhoods in a context-aware and modality-aware manner, something that sequential or flat representations cannot achieve effectively. While this study focuses on semantic and geographic signals, our methodology is not limited to these two modalities; without loss of generality, it can be extended to incorporate additional modalities such as temporal, sentiment, or even visual information, provided they can be represented in a graph-compatible format. We explore two complementary graph construction strategies. In the MonoGraph formulation, we define a unified graph $\mathcal{G}_{u}=(\mathcal{V},\mathcal{E}_{u})$, where edge weights integrate both modalities:
$w_{ij}=\mu\cdot\cos(\mathbf{h}_{i},\mathbf{h}_{j})+\theta\cdot\frac{1}{1+d(\mathbf{l}_{i},\mathbf{l}_{j})},$
where $h_{n}$ are the semantic embeddings and $l_{n}$ are the geographic coordinates, and $\mu$ and $\theta$ control the relative contributions of semantic and geographic similarities, respectively. This fused graph is then processed by a single Graph Convolutional Neural Network (GCN) encoder that learns joint representations capturing both semantic and spatial proximity. In the MultiGraph formulation, we construct two independent graphs over the same node set $\mathcal{V}$: a semantic graph $\mathcal{G}_{s}=(\mathcal{V},\mathcal{E}_{s})$, with edge weights:
$w_{ij}^{(s)}=\cos(\mathbf{h}_{i},\mathbf{h}_{j}),$
and a geographic graph $\mathcal{G}_{g}=(\mathcal{V},\mathcal{E}_{g})$, with edge weights:
$w_{ij}^{(g)}=\frac{1}{1+d_{\mathrm{hav}}(\mathbf{p}_{i},\mathbf{p}_{j})},$
where $\mathbf{p}_{i}=(\phi_{i},\lambda_{i})$ denotes the latitude–longitude coordinates of node $i$ in degrees under EPSG:4326, and $d_{\mathrm{hav}}$ is the great-circle (Haversine) distance given by:
@@
where $R$ is the Earth’s radius (e.g., $R\approx 6371\,\mathrm{km}$), and all angular quantities $\phi$ and $\lambda$ are expressed in radians before computation. Similarly, this graph is processed with a dedicated GCN network as well.
Figure 1: Layout of the two models: MonoGraph (a), and MultiGraph (b)

### v2

@@ 1 Introduction
Social media platforms continuously generate vast volumes of user-generated content, including posts, comments, images, videos, and interaction signals such as likes and shares. This content is rich, diverse, and real-time, offering a valuable resource for understanding how people communicate, express opinions, form communities, and respond to events. The widespread adoption of smartphones equipped with location-tracking capabilities has added an important additional dimension to this landscape: geographic context. By enabling users to share their locations, these devices have facilitated the integration of geospatial information into online content, significantly enhancing the contextual depth of social media analysis.
A growing body of research has demonstrated the value of geo-referenced social media data across a wide range of applications, including disaster response [goodchild2010crowdsourcing], public health monitoring [Signorini2011], and crisis management. In particular, location-aware content from microblogging platforms has been leveraged to study and monitor phenomena such as earthquakes [Havas2021Portability], floods [Huang2018Flood], refugee movements [Havas2021Refugee], and disease outbreaks [Stolerman2023Disease]. These studies highlight the potential of combining textual and spatial signals to gain timely and actionable insights from social media streams.
Much of the existing social media analysis literature has focused on semantic topic modeling as a fundamental technique for uncovering latent thematic structures within large collections of user-generated text [Resch2018, Hannigan2019]. While effective in capturing semantic regularities, these approaches typically treat text in isolation and rarely integrate additional modalities in a principled manner. Early efforts toward multimodal geo-social analysis often relied on sequential or modular workflows, in which each modality (e.g., text, location, time) was processed independently before being combined at a later stage [yin2011geographical]. Such designs limit the ability to capture complex cross-modal interactions and often introduce brittle dependencies between processing stages, making results sensitive to preprocessing choices and execution order.
More recent approaches have advanced toward integrating multiple modalities, including temporal signals, geographic space, semantic topics, and sentiment [gonzalez2024neural, hanny2025]. However, these methods still depend on fragmented pipelines with multiple isolated components, falling short of truly end-to-end multimodal learning. The inherent heterogeneity of input modalities further complicates this problem: text is symbolic and sequential, while geographic data is continuous and spatial. As a result, feature representations across modalities are often misaligned, leading to inconsistencies during fusion and reduced interpretability. The absence of a unified learning framework constrains cross-modal reasoning and limits the model’s ability to jointly exploit semantic and spatial structure.
In this paper, we propose a unified multimodal graph-based learning framework that jointly embeds semantic and geographic information into a shared representation space using end-to-end training. Our methodology constructs parallel semantic and geographic graphs over social media posts and learns representations that are both semantically coherent and spatially interpretable. Nodes in the graph correspond to post-level embeddings, while edges encode relational signals such as semantic similarity and geographic proximity. This formulation enables the model to reason over multimodal interactions in a structured and context-aware manner, rather than relying on post hoc fusion of independently learned features.
At the core of our approach is an unsupervised graph neural network that aggregates information from each node’s neighborhood through a weighted message-passing mechanism. A key innovation of the proposed framework is the use of a composite loss function inspired by reference-less clustering evaluation principles. The loss integrates three complementary objectives: a contrastive loss that preserves local semantic neighborhoods by pulling similar nodes closer while pushing dissimilar ones apart; a coherence loss that promotes compact and well-separated clusters; and an alignment loss that stabilizes training by encouraging consistency between node embeddings and their assigned cluster centroids. By embedding these objectives directly into the learning process, the model discovers topic clusters that are simultaneously semantically meaningful and geographically grounded.
Our approach is motivated by applications in disaster management, where geo-social media provides rich, fine-grained information from affected communities during emergencies. Through a case study of the 2017 Hurricane Harvey event in the United States, we demonstrate that the proposed framework produces geographically localized and semantically coherent topic clusters that closely align with observed disaster impacts. While this case study illustrates the practical utility of the model in crisis scenarios, the framework itself is application-agnostic and can be readily adapted to other multimodal data integration tasks across domains such as public health, urban analytics, and social sensing.
In summary, this work makes the following contributions:
@@
  • We introduce a unified multimodal graph-based framework that integrates semantic and geospatial information into a shared representation space for unsupervised topic discovery.
  • We propose two unsupervised architectures, MonoGraph and MultiGraph, that flexibly support different levels of multimodal interaction while remaining domain-agnostic.
@@
  • We develop a composite loss function combining contrastive, coherence, and alignment objectives to generate clusters that are both semantically coherent and spatially interpretable.
  • We validate the effectiveness of the proposed approach through a real-world disaster case study, demonstrating improved topic quality, spatial compactness, and interpretability compared to existing baselines.
@@ 2.1 Topic Modeling
Topic modeling is a widely used technique for discovering latent semantic topics within large document collections, providing a structured thematic summary by assigning one or more topic labels to each document [hoyle2022content]. Among the foundational methods, Latent Dirichlet Allocation (LDA) [blei2003latent] introduced a fully generative probabilistic framework in which each document is modeled as a mixture over latent topics, and each topic is represented as a multinomial distribution over words. The LDA framework has been extended to incorporate additional dimensions such as sentiment [Lin2009Joint], temporal dynamics [Blei2006Dynamic], and user-level information in social media [zhao2011comparing], broadening the applicability of topic models across domains. However, LDA and its variants fundamentally rely on the bag-of-words assumption, which ignores word order and semantic relationships between words, limiting their effectiveness in handling short, noisy, or highly contextual texts such as tweets and social media posts [hong2010empirical]. To overcome these limitations, embedding-based topic modeling techniques have recently emerged. Dieng et al. [dieng2020topic] introduced the Embedded Topic Model (ETM), which integrates pre-trained word and sentence embeddings with LDA’s generative framework to better capture semantic relationships. Sia et al. [sia2020neural] further demonstrated that clustering high-dimensional semantic embedding vectors, derived from neural language models, can serve as a powerful alternative strategy, achieving performance comparable to traditional topic models.
Building on these embedding-based approaches, Angelov [angelov2020top2vec] proposed Top2Vec, which leverages document embeddings and clustering algorithms to identify topics without requiring the number of topics to be pre-specified. This method utilizes techniques like Uniform Manifold Approximation and Projection (UMAP) and Hierarchical Density-Based Spatial Clustering of Applications with Noise (HDBSCAN) for dimensionality reduction and flexible clustering. Grootendorst [grootendorst2022bertopic] further advanced this line of work with BERTopic, which combines transformer-based contextual embeddings with refined clustering strategies, consistently outperforming traditional LDA and Top2Vec in both coherence and interpretability. Complementary empirical evidence from related domains further highlights the limitations of purely embedding-driven approaches. In particular, studies of pre-trained language models in simple knowledge graph question answering (KGQA) show that, while contextual embeddings capture strong lexical and semantic regularities, they struggle to reliably encode explicit relational structure without dedicated graph-based inductive bias [hu2023empirical]. These findings suggest that high-quality semantic embeddings alone are often insufficient for tasks requiring structured relational reasoning, motivating approaches that integrate neural language representations with explicit graph modeling.
@@ 2.2 Spatio-semantic Topic Modeling
As location-aware social media platforms proliferate, spatially grounded topic modeling has gained increasing attention. This line of work focuses on uncovering geographically coherent semantic patterns by integrating geospatial signals into the semantic topic modeling process. Early efforts extended classical models like LDA with spatial priors. For example, Spatial LDA (sLDA) [Yuan2013Who] and Geo-aware Topic Models [Hong2012Discovering] include location-based information as latent variables or observed metadata, enabling the discovery of region-specific topics. Similar models have been applied to a variety of domains, such as understanding natural disasters [Resch2018], flood mapping [Huang2018Flood], and epidemic monitoring [Stolerman2023Disease]. However, most spatial topic models treat location and text independently or as loosely coupled variables, which often leads to weak cross-modal alignment. In addition, spatial topic models are frequently built upon modular architectures, making them difficult to scale or generalize. More recent multimodal approaches attempt to combine geolocation, text, and other modalities, often through hybrid processing pipelines [gonzalez2024neural] [hanny2025]. While these methods capture richer context, they extract each modality feature in isolation and then merge features post hoc, limiting their ability to jointly reason across modalities. The absence of a unified representation space, together with the gap between symbolic textual data and continuous geospatial representations, further introduces alignment and interpretability challenges in downstream clustering or topic interpretation.
@@ 2.3 Graph-based Topic Modeling
Graph-based topic modeling has emerged as an effective alternative to traditional probabilistic models by explicitly exploiting relational structure to improve semantic coherence, particularly for short and sparse texts. Early approaches construct word co-occurrence graphs from text corpora and apply community detection algorithms to identify topical clusters. For instance, Wang et al. [wang2015community] demonstrated that community detection methods such as Louvain can outperform Latent Dirichlet Allocation (LDA) on short text datasets. Similarly, Zhao et al. [zhao2018graph] proposed the Graph-based Short Text Topic Model (GSTM), which leverages word co-occurrence graphs to alleviate data sparsity and improve topic coherence.
Subsequent work has sought to integrate graph structure more deeply into topic modeling. Newman and Girvan [newman2006finding] established a theoretical connection between community detection in bipartite document–word networks and topic model inference, highlighting the natural synergy between network science and probabilistic topic discovery. Extensions of LDA incorporating graph structure, such as Graph-LDA [doshi2015], embed word similarity relations directly into the generative framework, yielding more semantically coherent topics. More recently, Graph Topic Models (GTMs) have combined Graph Neural Networks (GNNs) with document–word heterogeneous graphs to jointly learn semantic representations and relational structure [zhou2020]. While these models benefit from relational inductive biases, they often struggle to capture deeper or long-range semantic dependencies and remain limited in modeling more abstract or nuanced topics.
Embedding-based graph clustering approaches further extend graph-based topic modeling by operating directly on learned sentence or document representations. Graph2Topic [graph2topic2023], for example, applies graph clustering to embedding-based representations, improving topic interpretability and quality without requiring a predefined number of topics. This line of work demonstrates the practical advantages of combining graph structure with embedding-based semantic representations for topic discovery. In parallel, recent advances in generic graph representation learning have introduced self-supervised contrastive objectives for learning node or graph embeddings without labels. GraphCL [you2020graphcl] performs graph-level contrastive learning by contrasting different augmented views of the same graph, while MVGRL [hassani2020] learns node representations by contrasting multiple structural views of a graph. Although these methods achieve strong performance on generic graph learning benchmarks, they are not designed for topic discovery and do not explicitly model semantic clustering, multimodal fusion, or geographically grounded objectives.
Beyond topic modeling, heterogeneous network embedding methods have been widely studied in domains such as recommendation systems, where multiple node and edge types are jointly modeled within a unified graph representation [shi2016heterogeneous, zhang2018recommendation]. These methods learn relational embeddings from heterogeneous graphs but are typically optimized for supervised or semi-supervised prediction objectives, rather than unsupervised topic discovery. As a result, they do not explicitly optimize for cluster-level semantic coherence, spatial grounding, or interpretability, which are central to geo-social topic modeling.
Across topic modeling, spatio-semantic extensions, and graph-based representation learning, prior work has largely addressed semantic structure, spatial information, and relational modeling in isolation. Classical and embedding-based topic models focus on semantic similarity, while spatially grounded topic models typically introduce geographic information as auxiliary metadata or enforce it as a post hoc constraint on semantically derived clusters [Yuan2013Who, Hong2012Discovering]. In parallel, recent graph-based multimodal fusion and contrastive learning methods emphasize learning robust relational embeddings, but are generally not designed for clustering-oriented topic discovery. Consequently, semantic structure, geographic proximity, and clustering objectives are rarely optimized jointly within a single end-to-end framework. Our framework addresses this gap by explicitly modeling semantic and geographic relations as parallel graph structures and integrating them within a unified learning objective tailored for unsupervised topic discovery, yielding clusters that are both semantically coherent and geographically grounded.
@@ 3 Methodology
Our approach leverages a collection of geotagged social media posts, where each post (node) $i$ is represented by a semantic embedding $\mathbf{h}_{i}$, derived from its textual content, and a geographic coordinate $\mathbf{l}_{i}$. Semantic representations are obtained using Sentence-BERT (SBERT), which fine-tunes transformer-based language models in a Siamese architecture to directly optimize sentence-level similarity. SBERT has been shown to effectively capture semantic proximity in short, informal texts such as tweets [Reimers2019]. To enable joint reasoning over heterogeneous data types, all modalities are embedded into graph-based representations that serve as a common computational substrate. As already mentioned, graphs provide a flexible framework for modeling local neighborhood structure and relational dependencies induced by semantic similarity and geographic proximity. While this study focuses on semantic and geographic signals, the methodology is formulated at the graph level and is therefore modality-agnostic, assuming additional information sources can be easyly encoded as node features or relational graphs. We explore two complementary graph construction strategies: a unified MonoGraph formulation and a decoupled MultiGraph formulation.
Figure 1: Layout of the two models: MonoGraph (a), and MultiGraph (b)
In the MonoGraph formulation, we construct a single heterogeneous graph $\mathcal{G}_{u}=(\mathcal{V},\mathcal{E}_{u})$ over social media posts, where each node $i\in\mathcal{V}$ corresponds to a post represented by a semantic embedding $\mathbf{h}_{i}$ and geographic coordinates $\mathbf{p}_{i}=(\phi_{i},\lambda_{i})$. The graph captures semantic and geographic information through two relation types defined over a shared node set. Semantic similarity between nodes is measured using cosine similarity, defining a semantic edge type with weights
$w_{ij}^{(s)}=\cos(\mathbf{h}_{i},\mathbf{h}_{j}).$
Geographic proximity is measured using the Haversine distance
@@
where $R\approx 6371\,\mathrm{km}$ denotes the Earth’s radius. Geographic proximity defines a geographic edge type with weights
$w_{ij}^{(g)}=\frac{1}{1+d_{\mathrm{hav}}(\mathbf{p}_{i},\mathbf{p}_{j})}.$
For each node $i$, a $k$-nearest-neighbor set is determined based on semantic similarity. Both semantic and geographic edge types are instantiated over this shared neighborhood, without introducing explicit modality weighting. Graph sparsity is controlled solely through $k$-nearest neighbor selection. The resulting heterogeneous graph is processed by a heterogeneous GCN encoder composed of relation-specific graph convolution operators. At each layer, messages from different relations are aggregated via summation, enabling joint propagation of semantic similarity and spatial proximity within a unified message-passing framework.
In the MultiGraph formulation, semantic and geographic relations are instantiated as two independent graphs over the same node set $\mathcal{V}$, allowing modality-specific neighborhood structures to be modeled separately. The semantic graph $\mathcal{G}_{s}=(\mathcal{V},\mathcal{E}_{s})$ connects each node to its top-$k$ nearest neighbors based on cosine similarity of semantic embeddings, using the semantic edge weights $w_{ij}^{(s)}$ defined above. Graph sparsity is controlled solely through $k$-nearest-neighbor selection, with no global similarity threshold. In parallel, the geographic graph $\mathcal{G}_{g}=(\mathcal{V},\mathcal{E}_{g})$ is constructed by connecting each node to its $k$ geographically nearest neighbors based on great-circle distance, with geographic edge weights $w_{ij}^{(g)}$ computed using the Haversine distance defined above. No additional thresholds or modality-specific reweighting are introduced. Each graph is processed by a dedicated GCN encoder, enabling semantic and geographic neighborhood information to be learned independently prior to fusion.